### v1

# SloPalSpeech: A 2,800-Hour Slovak Speech Corpus from Parliamentary Data

Abstract
Automatic Speech Recognition (ASR) for low-resource languages like Slovak is hindered by the scarcity of training data. To address this, we introduce SloPalSpeech, a new, large-scale Slovak ASR dataset containing 2,806 hours of speech from parliamentary proceedings. We developed a robust processing pipeline to align and segment long-form recordings into clean, 30-second audio-transcript pairs suitable for model training. We use this dataset to fine-tune several OpenAI Whisper models (small, medium, large-v3, and large-v3-turbo), achieving significant Word Error Rate (WER) reductions on standard Slovak benchmarks like Common Voice and FLEURS. For instance, the fine-tuned Whisper-small model’s WER dropped by up to 70%, approaching the baseline performance of the much larger Whisper-large-v3 model. To foster future research in low-resource speech recognition, we publicly release the complete SloPalSpeech dataset, the fully segmented transcripts (60 million words), and all our fine-tuned models.
Keywords: Slovak ASR, Speech recognition, Parliamentary data, Whisper fine-tuning
[…]
SloPalSpeech: A 2,800-Hour Slovak Speech Corpus from Parliamentary Data
Erik Božík1,4   Marek Šuppa2,3,4
1VÚB Banka, Bratislava, Slovakia
2Faculty of Mathematics, Physics and Informatics, Comenius University in Bratislava
3Cisco Systems, Inc. 4NaiveNeuron
Correspondence: ebozik@vub.sk
Abstract content
1.  Introduction
Automatic Speech Recognition (ASR) has improved significantly with large multilingual datasets and models such as Whisper Radford et al. (2023). However, support for Slovak remains limited. To address this gap, we explore publicly available parliamentary proceedings as a large-scale source of Slovak speech data. We collect and process thousands of hours of recordings and transcripts to create an ASR dataset and fine-tune multiple Whisper models.
Public ASR benchmarks, such as the Hugging Face Open ASR Leaderboard Srivastav et al. (2023), reveal substantial variation among top-performing models from diverse contributors. However, support for low-resource languages like Slovak remains limited. Many leading models either lack Slovak support entirely or are proprietary, restricting their usability for research. From the models shown in Table 1, the first openly available model with Slovak support is CrisperWhisper Zusag et al. (2024), a modified version of OpenAI’s Whisper large-v3.
Model Avg WER
nvidia/canary-gwen-2.5b 5.63
ibm-granite/granite-speech-3.3-8b 5.85
nvidia/parakeet-tdt-0.6b-v2 6.05
microsoft/Phi-4-multimodal-instruct 6.14
nvidia/canary-1b-flash 6.35
kyutai/stt-2.6b-en 6.40
nvidia/canary-1b 6.50
nyrahealth/CrisperWhisper 6.67
ibm-granite/granite-speech-3.3-2b 6.86
elevenlabs/scribe_v1 6.88
speechmatics/enhanced 6.91
Table 1: Average WER (%) of top-rated models in Open ASR leaderboard at the time of writing
Models like OpenAI’s Whisper Radford et al. (2023) rely on large, diverse datasets, yet public resources remain scarce for many languages, leading to weaker performance. As shown in Table 2 and Table 3, Whisper achieves much lower WERs on high-resource languages (e.g., English, Russian, German; thousands of training hours) than on low-resource ones (e.g., Slovak, Croatian, Lithuanian; $\sim$100 training hours). Large models can reach 3–5% WER in high-resource settings, whereas low-resource languages often struggle to go below 10%, underscoring the need for more Slovak training data to close this gap.
Model Slovak Croatian Lithuanian
Whisper tiny 77.2 79.0 98.5
Whisper small 33.3 33.4 65.6
Whisper medium 17.3 19.3 41.1
Whisper large‑v2 11.7 13.4 28.1
Hours 90 91 67
Table 2: Whisper WERs on FLEURS on selected low-resource languages
Model English Russian German
Whisper tiny 12.4 31.1 27.8
Whisper small 6.1 11.4 10.2
Whisper medium 4.4 7.2 6.5
Whisper large‑v2 4.2 5.6 4.5
Hours 438k 9.8k 13.3k
Table 3: Whisper WERs on FLEURS on selected high-resource languages
2.  Objectives
The objective of this work is to address the challenges faced by low-resource languages, as outlined in Section 1, with a particular focus on the Slovak language. Our primary goal is to explore existing resources of Slovak parliamentary speech and construct a large-scale dataset suitable for training speech recognition models. This dataset will consist of audio clips, each no longer than 30 seconds, paired with corresponding textual transcriptions.
Publicly available resources rarely exist in this format, making it likely that preprocessing will be required. As a result, some data may be lost during this process. Therefore, we also aim to investigate the possibility of obtaining fully transcribed textual data, even when no matching audio is available, to maximize the utility of available resources.
The 30-second clip length is chosen in alignment with the Whisper family of models, which we intend to fine-tune using the curated dataset. Our goal is to develop a set of models optimized for Slovak speech recognition. These will include fine-tuned versions of models from Whisper-small up to Whisper-large-v3 (excluding versions v1 and v2). By training models across different sizes, we aim to assess the impact of our dataset on model performance and language-specific accuracy.
3.  Dataset
Several open-source datasets exist for Slovak, such as CommonVoice Ardila et al. (2019), FLEURS Conneau et al. (2023), and VoxPopuli Wang et al. (2021), but together they offer only about 100 hours of data which is far below the scale required for high-performance ASR. We focus on Slovak parliamentary hearings as a promising resource to bridge this gap, potentially providing thousands of hours of training data and far surpassing existing datasets like those mentioned. These hearings are fully public and regularly transcribed, making them a strong candidate for ASR dataset creation. Audio recordings and their corresponding transcripts are publicly available on various government websites, with coverage dating back to the establishment of the Slovak Republic. This long-term availability presents a valuable opportunity to construct a dataset that qualifies as large-scale. To create this, we must first assess their original availability.
3.1.  MediaPortál
Parliamentary recordings and transcripts are available across multiple platforms. While older sessions are accessible through a legacy portal with inconsistent coverage, MediaPortál NRSR has emerged as the primary and most comprehensive source for more recent election terms. It provides a centralized interface for browsing and accessing audio recordings of parliamentary hearings, making it the most suitable foundation for building our ASR dataset.
Hearings in MediaPortál follow a structure similar to the one shown in Figure 1. From the layout, users can navigate through the entire recording by simply clicking on the speaker segments. However, we observed that these segments were often inaccurate. Due to these inconsistencies, we concluded that, in order to construct a dataset based on our predefined structure, the data would need to be re-aligned. As a result, we abandoned the idea of directly using the provided segments and instead focused on collecting the full-text transcript. While it is possible to generate the transcript by concatenating all individual segments, we opted to use the official administrative version of the transcript, which is readily available in full-text form on The Joint Czech and Slovak Digital Parliamentary Library National Council of the Slovak Republic and Parliament of the Czech Republic (2002).
Figure 1: Sample page from MediaPortál
3.2.  The Joint Czech and Slovak Digital Parliamentary Library
This initiative, established in 2002 in Slovakia, aims to digitize parliamentary documents. Since verbatim transcripts of parliamentary hearings fall within this category, these documents can be easily filtered and accessed in digital form through the portal. The platform features a tree-view interface, which offers an intuitive way to filter transcripts for each session, as illustrated in Figure 2. Under each session, multiple links are provided - each corresponding to the verbatim transcript of a specific meeting day, available in DOCX format.
Our examination of transcripts across different years revealed that documents which cover all the session recordings available in MediaPortál follow a consistent formatting structure, as shown in Figure 3. The layout is organized around speaker announcements, which are always presented in bold. These announcements follow a standardized format: surname, first name, and the speaker’s role at the time, each separated by a comma. Bold text is also used beyond speaker announcements - for instance, in headers that convey important information such as the date and time of the session. Between each speaker announcement appears the actual verbatim transcript of their speech.
Figure 2: Example layout from the Digital parliamentary library
Figure 3: Example transcript in .docx format
$3.$  Collection
We collected all available audio recordings from MediaPortál. Since the frontend download functionality was not working, we extracted the URLs of the HLS streams and used FFmpeg FFmpeg Developers (2025) to download the recordings. Due to storage limitations, we opted to extract only the audio in MP3 format. This resulted in a total of 4512.5 hours of audio.
For the transcripts, we directly downloaded the available DOCX documents. Some documents had to be discarded because the webpage redirected to an HTML page instead of a valid DOCX file. In total, we successfully collected 2111 transcript files.
To match each audio recording with the corresponding transcript, we extracted both the session number and the exact date. This was necessary because a single session number can span multiple days. As a result, the session number alone was not sufficient to produce an accurate match. Therefore, the date was also required to uniquely identify each recording. Using this method, we were able to match 1136 recordings, covering a total of 4098 hours of audio.
3.4.  Parsing
The first step in effectively parsing DOCX files was to convert them into a format that is easier to process. For this purpose, we used Apache Tika (specifically the Tika Server) Apache Software Foundation (2025) which enabled conversion to XHTML. This transformation allowed us to perform HTML-like parsing, as long as XHTML-specific features were disregarded. As a result of the conversion, some formatting elements were preserved as tags, most notably bold text enclosed in <b></b> tags.
To filter out all non-verbatim content, we needed to segment the transcript based on speaker annotations. However, due to inconsistencies caused by human error (missing commas, inconsistent formatting of names) or edge cases like three-part names, we were unable to rely on a simple split based on commas or regular expressions. Therefore, we adopted a heuristic-based approach using the following conditions:
  1. 1.
    The text was bold.
  2. 2.
    It contained at least one and no more than three names.
  3. 3.
    It did not exceed 15 words.
To support this logic, we scraped a complete list of National Council members from all electoral terms using the official parliamentary website NRSR . From this data, we compiled a comprehensive set of known names, including both first and last names for each member. Each time a bold line was encountered, we checked it against this set. If the required number of name matches was found, the line was classified as a speaker annotation. We also applied a word limit, as we observed that this conditions alone matched some text that was not a speaker annotation but a spoken text. After applying these steps, we achieved a structured format as shown in Listing 4.
Additionally, we needed to filter out all text enclosed in parentheses, as such content typically represents transcriber’s notes. Specifically, we removed all text enclosed in both () and []. To handle cases where parentheses were not properly closed, we implemented a pattern that treated a bracket as the start of the note and a period as the end.
"speaker": "Laššáková, Jana, podpredsedníčka NR SR",
"transcript": "Príjemné, dobré ráno, vážené panie poslankyne, páni poslanci…"
"speaker": kripek, Branislav, poslanec NR SR",
"transcript": akujem pekne, pani predsedajúca. Áno, preto som sa chcel vydýchať…"
Figure 4: JSON example of parsed transcript
3.5.  Validity
To ensure the transcripts matched the audio, we randomly sampled five sessions and created a test set. We then evaluated several Whisper models on the audio and measured the WER against the provided normalized transcripts. [^1] We used the WhisperX framework Bain et al. (2023) for its ability to process long-form audio, acknowledging that results may differ slightly with alternative splitting methods or without the CTranslate2 Whisper implementation. Our goal was to validate the transcripts rather than precisely evaluate model performance. Initial WER values exceeded 1.0, indicating a major mismatch. This led us to discover that the audio only included the forenoon part of sessions, while the transcripts covered both forenoon and afternoon. After removing the afternoon parts from the transcripts, we obtained WERs shown in Table 4, which closely match those reported by the Whisper authors. We therefore consider the transcripts valid for the forenoon parts.
Model Word Error Rate
openai/whisper-large-v3 12.9%
openai/whisper-medium 21.7%
openai/whisper-small 36.1%
openai/whisper-tiny 77.3%
Table 4: WER of different Whisper models (their CTranslate2 implementation) on parliamentary transcripts
4.  Alignment and segmentation
The collected data consisted of transcripts and audio recordings of varying lengths, ranging from a few seconds up to 25 hours, with an average duration of 3.6 hours. This diverged significantly from our target structure, which limits segments to a maximum of 30 seconds. As noted in Section 3.1, this made both alignment and segmentation necessary to meet our required dataset structure.
Although forced alignment is a common method for generating timestamps, it was not suitable in our case. The transcripts frequently included more content than was actually spoken, and the long duration of the recordings exceeded the effective input range of most forced alignment models, which typically perform better on shorter segments. To overcome these limitations, we developed an alternative approach based on generating a reference transcript from the audio and aligning it with the collected transcript, which we treated as the ground truth.
4.1.  Reference Transcript
We first generated a reference transcript with word-level timestamps using the WhisperX framework, which combines Whisper’s transcription capabilities with external alignment models Bain et al. (2023). While Whisper provides reasonably accurate results (with a WER of 12.9% for the largest model, as shown in Table 4), the output still contained typical inaccuracies. However, since it outputs transcripts in 30-second chunks, it enabled effective forced alignment afterward without the limitations previously mentioned.
4.2.  Anchor-Based Timestamping
To align the reference transcript with the ground truth, we identified “anchors” - words that appear in both versions and can serve as reliable reference points. We assumed that some words in the reference were correctly transcribed and could be matched to their counterparts in the ground truth. Each reference word was matched to candidate ground truth words within a Levenshtein distance of 1 Levenshtein (1966), provided the match did not jump backward relative to the last aligned word and stayed within a 50-word forward window. If no valid candidates were found, the word was skipped.
Among all candidates, we selected the best match by comparing the context: specifically, the four preceding and four following words. Each candidate received a score based on the number of full-text matches in this window, and the one with the highest score was chosen as the anchor. To further improve accuracy, we applied several heuristics, such as skipping words shorter than three characters and requiring a minimum match score of 3 for a word to qualify as an anchor.
This process yielded a list of ground truth words with corresponding timestamps. Because each match was constrained to move forward, the resulting timestamps formed a strictly increasing sequence, which effectively eliminated alignment anomalies.
4.3.  Segment Construction from Anchors
The anchoring method proved reasonably effective, as most sessions contained a sufficient number of matched words. To construct segments suitable for our ASR dataset, we aimed to group transcript and audio pairs that were no longer than 30 seconds.
We iterated through the list of anchors, starting from each one and scanning forward until we found the next anchor with a time difference greater than 28 seconds. We used a 2-second buffer to create a soft threshold. All ground truth words between the starting anchor and this later anchor formed a candidate segment, and the corresponding audio slice was extracted using their timestamps.
This approach yielded a useful number of segments across most sessions. However, in cases with sparse anchors - caused by long pauses or local misalignments - a segment was still constructed, but its validity required further assessment.
$4.$  Final Selection
Although the algorithm itself was deterministic, it operated on Whisper-generated transcripts, which are inherently variable and sometimes unpredictable. This occasionally led to anomalies or undesired behavior in specific contexts. To address this, we applied a post-processing step where the Whisper model was reapplied to each segmented audio chunk shorter than 30 seconds [^2] This time without WhisperX since Whisper could be applied directly - enabling faster inference, generating updated transcripts and subsequently calculating the WER. The resulting distribution is shown in Figure 5. While most segments had a WER below 50%, a notable portion exhibited significantly higher values. We analyzed different WER intervals and found that lower rates typically indicated well-aligned and correctly segmented audio, whereas higher WERs were often associated with missegmentations or transcription errors. This trend was confirmed through manual inspection across the range of values.
To remove problematic segments while preserving as much usable content as possible, we empirically selected a WER threshold of 40%. Applying this filter reduced the dataset from 2977.73 to 2806 hours, eliminating 171.73 hours of low-quality data. In terms of segments, the count decreased from 427,276 to 402,966 - a reduction of 24,310 segments. This threshold effectively retained the majority of the data while substantially improving its overall quality.
Figure 5: Distribution of WER values on segmented audio
5.  Fine-tuning
With the dataset prepared, we proceeded to fine-tune multiple Whisper models to then be able to assess the impacts of parliamentary speech data on Slovak ASR performance. This section outlines our experimental setup and the fine-tuning strategies adopted for different model sizes, including the adaptations required for the largest variant.
5.1.  General Setup
We utilized multiple NVIDIA A10 GPUs to accelerate training. Given that Whisper has been widely adopted since its 2022/2023 release, we implemented the fine-tuning pipeline using Hugging Face Transformers, which simplifies much of the low-level implementation. Our fine-tuning strategy involved training for a maximum of three epochs with early stopping, based on evaluation performance. This setup was informed by prior research on transformer models such as Devlin et al. (2018), as well as the original Whisper pre-training protocol Radford et al. (2023), which also relied on a small number of epochs.
Although we had over 2,800 hours of training data, its formal and domain-specific nature raised concerns about catastrophic forgetting. The model could adapt well to parliamentary speech but lose generalization capability for broader Slovak. To assess this, we monitored performance not only on a held-out parliamentary test set but also on subsets of FLEURS and Common Voice. For both of these datasets we produced a subset of 300 samples. These additional benchmarks provided insight into how well the model retained general Slovak transcription capabilities across more diverse contexts. The parliamentary test set was of similar size.
5.2.  Whisper Small, Medium, Turbo
Fine-tuning these models was feasible on a single GPU due to their manageable size. We based our code on a blog post by Gandhi (2022) and adopted most of the original hyperparameters, with the exception of the learning rate. While many works default to a fixed value of $1\times 10^{-5}$, we followed a recommendation by Lodagala (2023) of using a learning rate 40× smaller than that used during pretraining. In our experiments, this adjustment led to more stable convergence across most of the model sizes.
We applied early stopping between the 2nd and 3rd epoch after observing evaluation plateaus on all of the test sets. None of the models showed signs of overfitting or catastrophic forgetting. All achieved significant improvements on the parliamentary test set, and importantly, also generalized well to broader Slovak speech - gradually yielding notable WER reductions on both the FLEURS and Common Voice subsets during training.
5.3.  Whisper Large v3
Fine-tuning Whisper Large v3 required adjustments beyond those used for smaller models, mainly due to its size exceeding single-GPU memory limits. While official details on its training are unavailable, it is known to have been trained on millions of hours of data and to use a differently sized Mel spectrogram than earlier versions. These differences initially led to configuration issues and overfitting, which we had to address.
5.3.1.  Multi-GPU setup
We addressed the GPU limitations by using Fully Sharded Data Parallel (FSDP) Zhao et al. (2023). Although Hugging Face Transformers Wolf et al. (2020) officially support FSDP and multi-GPU training via Accelerate Gugger et al. (2022), we faced persistent compatibility issues - different library versions produced inconsistent runtime errors with the same code. Training was eventually stabilized using version 4.52.4, but evaluation consistently failed due to runtime errors. As a workaround, we used three GPUs for training and a fourth GPU to run inference at selected checkpoints. This differed from the single-GPU setup used for smaller models, where training and evaluation could run together in a single process.
We used a sharded state dict configuration, which produced FSDP checkpoints during training. These were merged into a single safetensors file, which we then converted back to its original format with split weights, matching the structure used in the original release.
5.3.2.  Overfitting
Starting with the same setup as previous models, we used a learning rate 40× smaller, as mentioned earlier. However, after the warmup stage, WER quickly increased - first on the parliamentary test set, then on the other evaluation sets. This indicated significant overfitting, which needed to be addressed to obtain a usable model. As noted in forum discussions Huggingface (2024), others reported encountering similar issues, though no definitive solution was identified. We reverted to the commonly reported settings, excluding LoRA adapters, as LoRA was not our objective. This included adjusting the learning rate to $1\times 10^{-5}$. After this we proceeded with hyperparameter experimentation. Introducing a weight decay of 0.01 (previously 0) enabled stable convergence in stages that had previously overfitted, reducing WER across all test sets. This became our final configuration for Whisper Large v3 with which we proceeded until an early stopping around 2 epochs.
6.  Benchmarks
To evaluate the impact of parliamentary speech fine-tuning on Slovak ASR performance, we conducted a comprehensive benchmarking of all produced models. While preliminary indicators were available from the training process, final assessments were carried out on the official test sets of Common Voice 21 and FLEURS. As with all benchmarking we included basic text normalization before producing the WER value.
6.1.  Improvements
Model CV21 FLEURS $\Delta$ (%)
Small-Base 58.4 36.1 32.7-32.7 / 25.5-25.5
Small-Tuned 25.7 10.6
Medium-Base 38.0 18.7 20.0-20.0 / 11.1-11.1
Medium-Tuned 18.0 7.6
Table 5: WER comparison of Whisper Small and Medium (Base vs fine-tuned). $\Delta$ indicates absolute WER reduction (%) for CV21 / FLEURS.
Model CV21 FLEURS $\Delta$ (%)
Turbo-v3 Base 31.7 10.7 18.5-18.5 / 4.3-4.3
Turbo-v3 Tuned 13.2 6.4
Large-v3 Base 20.8 9.2 9.2-9.2 / 3.7-3.7
Large-v3 Tuned 11.6 5.5
Table 6: WER comparison of Whisper Large-v3 variants (Base vs fine-tuned). $\Delta$ indicates absolute WER reduction (%) for CV21 / FLEURS.
The benchmarking outcomes, presented in Table 5 for Whisper Small and Medium and in Table 6 for Whisper Large-v3-Turbo and Large-v3, show WER decreases across all models, with the largest relative improvements observed in the smaller ones. For Whisper Small, fine-tuning reduced WER by approximately 65–70% compared to the baseline, approaching the base performance of Large-v3. This demonstrates that the fine-tuned dataset effectively adapted model weights, with diminishing marginal gains as model size increased. Manual inspection confirmed that fine-tuning improved both word-level accuracy and the generation of Slovak-like token compositions, even when predictions were partially incorrect. Whisper Medium showed similar but smaller improvements, still yielding substantial performance gains.
Despite Whisper Large-v3 being trained on millions of hours of data, our dataset further reduced WER by a few percentage points. At these low error rates, improvements were less apparent during manual inspection, though handling of rare Slovak words noticeably improved. Among all evaluated models, Whisper Large-v3-Turbo offered the best performance-to-size trade-off: with roughly 730 million fewer parameters than Large-v3 and only about 100 million more than Medium, it achieved accuracy within about 1% of Large-v3, making it a strong candidate for efficient inference.
7.  Results
We compiled a large-scale dataset of 2.8k hours of Slovak parliamentary hearings paired with transcripts, suitable for ASR model training. Due to the long-form nature of the recordings and transcripts, our alignment method could not cover the entire collection. Therefore, in addition to the ASR dataset, we release all speaker-segmented transcripts, comprising approximately 60 million words.
To validate the dataset and develop accurate Slovak ASR models for real-world applications, we fine-tuned four Whisper models of varying sizes. Evaluated on Common Voice 21 and FLEURS, all models achieved reduced WER, with particularly large relative improvements for smaller models and notable gains for larger ones.
8.  Conclusion
We presented a complete pipeline for creating a large-scale ASR dataset, from sourcing and collecting raw data to alignment, segmentation, and demonstrating its effectiveness in training real-world models. With 2.8k hours of speech, this work helps move Slovak ASR beyond low-resource status. We further show the strong utility of parliamentary speech for Slovak ASR training and provide a methodology for aligning long-form audio with extensive transcripts.
We publish all resources, including the ASR dataset (2.8k hours [^3] This dataset includes only the portion of audio successfully aligned with transcripts. The full set of segmented transcripts (60 million words) is a superset of the text in the ASR dataset.), the complete collection of segmented transcripts (60 million words), and the fine-tuned Whisper models (small, medium, large-v3-turbo, and large-v3).
Limitations
Whisper models are known to hallucinate due to their architecture. While the base model often repeated words like "Ďakujem," our fine-tuned models occasionally produced sequences such as "Pán poslanec," reflecting a bias toward parliamentary speech. We recommend adopting an approach similar to Faster Whisper SYSTRAN (2025), where compression ratio checks during inference can trigger a re-decode with adjusted temperature to mitigate hallucinations. Additionally, when encountering unknown words or phrases, the model may substitute them with parliamentary terms. For example, in one Common Voice sample with the transcript "ostrov Nihau" the model predicted "ostrov Mihál" where "Mihál" is a former politician’s surname. We assume that similar cases are likely to happen.
Since the fine-tuning process updated all model parameters, the model shifted strongly toward Slovak. While we did not quantify the effect, we observed degraded English transcription performance, suggesting a loss of multilingual capability.
9.  Future Work
Parliamentary proceedings are transcribed in many European countries, where several languages remain low-resource for ASR. With access to recordings and full transcripts, our alignment method could be scaled to other languages to boost ASR resources beyond Slovak.
10.  Acknowledgements
We thank VÚB Banka for their generous support and for providing the GPU resources necessary to accomplish this work.
11.  Bibliographical References
[…]
  • Apache Software Foundation (2025) Apache Software Foundation. 2025. Apache tika project wiki. https://cwiki.apache.org/confluence/display/TIKA/Home.
  • Ardila et al. (2019) Rosana Ardila, Megan Branson, Kelly Davis, Michael Henretty, Michael Kohler, Josh Meyer, Reuben Morais, Lindsay Saunders, Francis M Tyers, and Gregor Weber. 2019. Common voice: A massively-multilingual speech corpus. arXiv preprint arXiv:1912.06670.
  • Bain et al. (2023) Max Bain, Jaesung Huh, Tengda Han, and Andrew Zisserman. 2023. Whisperx: Time-accurate speech transcription of long-form audio. arXiv preprint arXiv:2303.00747.
  • Conneau et al. (2023) Alexis Conneau, Min Ma, Simran Khanuja, Yu Zhang, Vera Axelrod, Siddharth Dalmia, Jason Riesa, Clara Rivera, and Ankur Bapna. 2023. Fleurs: Few-shot learning evaluation of universal representations of speech. In 2022 IEEE Spoken Language Technology Workshop (SLT), pages 798–805. IEEE.
  • Devlin et al. (2018) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. CoRR, abs/1810.04805. ArXiv:1810.04805.
  • FFmpeg Developers (2025) FFmpeg Developers. 2025. Ffmpeg documentation.
  • Gandhi (2022) Sanchit Gandhi. 2022. Fine‑Tune Whisper For Multilingual ASR with Transformers.
  • Gugger et al. (2022) Sylvain Gugger, Lysandre Debut, Thomas Wolf, Philipp Schmid, Zachary Mueller, Sourab Mangrulkar, Marc Sun, and Benjamin Bossan. 2022. Accelerate: Training and inference at scale made simple, efficient and adaptable. https://github.com/huggingface/accelerate.
  • Huggingface (2024) Huggingface. 2024. Whisper large-v3 finetuning - forum post.
  • Levenshtein (1966) Vladimir I. Levenshtein. 1966. Binary codes capable of correcting deletions, insertions and reversals. Soviet Physics Doklady, 10(8):707–710. Originally published in Russian in Doklady Akademii Nauk SSSR (1965).
  • Lodagala (2023) Vasista Sai Lodagala. 2023. Fine-tuning and evaluating whisper models for automatic speech recognition.
  • National Council of the Slovak Republic and Parliament of the Czech Republic (2002) National Council of the Slovak Republic and Parliament of the Czech Republic. 2002. Joint czech and slovak digital parliamentary library. https://www.nrsr.sk/dl/.
  • (13) NRSR. Abecedný zoznam poslancov národnej rady slovenskej republiky. https://www.nrsr.sk/web/?sid=poslanci/zoznam_abc.
  • (14) NRSR. Mediaportál národnej rady slovenskej republiky. https://tv.nrsr.sk/.
  • Radford et al. (2023) Alec Radford, Jong Wook Kim, Tao Xu, Greg Brockman, Christine McLeavey, and Ilya Sutskever. 2023. Robust speech recognition via large-scale weak supervision. In International conference on machine learning, pages 28492–28518. PMLR.
  • Srivastav et al. (2023) Vaibhav Srivastav, Somshubra Majumdar, Nithin Koluguri, Adel Moumen, Sanchit Gandhi, et al. 2023. Open automatic speech recognition leaderboard.
  • SYSTRAN (2025) SYSTRAN. 2025. Systran/faster-whisper. GitHub repository.
  • Wang et al. (2021) Changhan Wang, Morgane Riviere, Ann Lee, Anne Wu, Chaitanya Talnikar, Daniel Haziza, Mary Williamson, Juan Pino, and Emmanuel Dupoux. 2021. Voxpopuli: A large-scale multilingual speech corpus for representation learning, semi-supervised learning and interpretation. arXiv preprint arXiv:2101.00390.
  • Wolf et al. (2020) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, Joe Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien Plu, Canwen Xu, Teven Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest, and Alexander M. Rush. 2020. Transformers: State-of-the-art natural language processing. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 38–45, Online. Association for Computational Linguistics.
  • Zhao et al. (2023) Yanli Zhao, Andrew Gu, Rohan Varma, Liang Luo, Chien-Chin Huang, Min Xu, Less Wright, Hamid Shojanazeri, Myle Ott, Sam Shleifer, et al. 2023. Pytorch fsdp: experiences on scaling fully sharded data parallel. arXiv preprint arXiv:2304.11277.
  • Zusag et al. (2024) Mario Zusag, Laurin Wagner, and Bernhad Thallinger. 2024. Crisperwhisper: Accurate timestamps on verbatim speech transcriptions. In Proc. Interspeech 2024, pages 1265–1269.